% Generated by IEEEtran.bst, version: 1.14 (2015/08/26)
\begin{thebibliography}{100}
\providecommand{\url}[1]{#1}
\csname url@samestyle\endcsname
\providecommand{\newblock}{\relax}
\providecommand{\bibinfo}[2]{#2}
\providecommand{\BIBentrySTDinterwordspacing}{\spaceskip=0pt\relax}
\providecommand{\BIBentryALTinterwordstretchfactor}{4}
\providecommand{\BIBentryALTinterwordspacing}{\spaceskip=\fontdimen2\font plus
\BIBentryALTinterwordstretchfactor\fontdimen3\font minus
  \fontdimen4\font\relax}
\providecommand{\BIBforeignlanguage}[2]{{%
\expandafter\ifx\csname l@#1\endcsname\relax
\typeout{** WARNING: IEEEtran.bst: No hyphenation pattern has been}%
\typeout{** loaded for the language `#1'. Using the pattern for}%
\typeout{** the default language instead.}%
\else
\language=\csname l@#1\endcsname
\fi
#2}}
\providecommand{\BIBdecl}{\relax}
\BIBdecl

\bibitem{Malikopoulos2022a}
A.~A. Malikopoulos, ``Separation of learning and control for cyber-physical
  systems,'' \emph{Automatica}, vol. 151, no. 110912, 2023.

\bibitem{Malikopoulos2011b}
------, \emph{{Real-Time, Self-Learning Identification and Stochastic Optimal
  Control of Advanced Powertrain Systems}}.\hskip 1em plus 0.5em minus
  0.4em\relax ProQuest, 2011.

\bibitem{Malikopoulos2007}
A.~A. Malikopoulos, P.~Y. Papalambros, and D.~N. Assanis, ``{A State-Space
  Representation Model and Learning Algorithm for Real-Time Decision-Making
  Under Uncertainty},'' in \emph{ASME 2007 International Mechanical Engineering
  Congress and Exposition}, 2007, pp. 575--584.

\bibitem{Malikopoulos2007a}
------, ``{A learning algorithm for optimal internal combustion engine
  calibration in real time},'' in \emph{ASME 2007 International Design
  Engineering Technical Conferences and Computers and Information in
  Engineering Conference}, 2007, pp. 91--100.

\bibitem{Malikopoulos2009b}
A.~A. Malikopoulos, ``{Convergence properties of a computational learning model
  for unknown Markov chains},'' \emph{J. Dyn. Sys., Meas., Control}, vol. 131,
  no.~4, pp. 041\,011--7, 2009.

\bibitem{Malikopoulos2009a}
A.~Malikopoulos, D.~Assanis, and P.~Papalambros, ``{Real-time self-learning
  optimization of diesel engine calibration},'' \emph{Journal of Engineering
  for Gas Turbines and Power}, vol. 131, no.~2, 2009.

\bibitem{Malikopoulos2006}
A.~Malikopoulos, Z.~Filipi, and D.~Assanis, ``{Simulation of an Integrated
  Starter Alternator (ISA) system for the HMMWV},'' \emph{SAE Technical
  Papers}, 2006.

\bibitem{Malikopoulos2011a}
A.~Malikopoulos, ``{Equilibrium control policies for Markov chains},'' in
  \emph{Proceedings of the 50th IEEE Conference on Decision and Control}, 2011,
  pp. 7093--7098.

\bibitem{Malikopoulos2015}
A.~Malikopoulos, V.~Maroulas, and J.~Xiong, ``{A multiobjective optimization
  framework for stochastic control of complex systems},'' in \emph{Proceedings
  of the American Control Conference}, 2015, pp. 4263--4268.

\bibitem{Malikopoulos2015b}
A.~Malikopoulos, ``{Centralized stochastic optimal control of complex
  systems},'' in \emph{2015 European Control Conference}, pp. 721--726.

\bibitem{Malikopoulos2008b}
A.~Malikopoulos, D.~Assanis, and P.~Papalambros, ``{Optimal engine calibration
  for individual driving styles},'' \emph{SAE Technical Papers}, 2008.

\bibitem{Malikopoulos2010a}
A.~Malikopoulos, ``{A rollout control algorithm for discrete-time stochastic
  systems},'' in \emph{ASME 2010 Dynamic Systems and Control Conference,
  DSCC2010}, vol.~2, 2010.

\bibitem{Malikopoulos2010}
A.~A. Malikopoulos, ``{A Rollout Control Algorithm for Discrete-Time Stochastic
  Systems},'' \emph{Development}, pp. 1--7, 2010.

\bibitem{Malikopoulos2009e}
------, ``{Convergence properties of a computational learning model for unknown
  markov chains},'' in \emph{2008 Proceedings of the ASME Dynamic Systems and
  Control Conference}, 2008, pp. 73--80.

\bibitem{Malikopoulos2010b}
A.~Malikopoulos, P.~Papalambros, and D.~Assanis, ``{Online identification and
  stochastic control for autonomous internal combustion engines},''
  \emph{Journal of Dynamic Systems, Measurement and Control, Transactions of
  the ASME}, vol. 132, no.~2, pp. 1--6, 2010.

\bibitem{Malikopoulos2009}
------, ``{A real-time computational learning model for sequential
  decision-making problems under uncertainty},'' \emph{Journal of Dynamic
  Systems, Measurement and Control, Transactions of the ASME}, vol. 131, no.~4,
  2009.

\bibitem{Malikopoulos2013d}
A.~A. Malikopoulos, ``{Method, control apparatus and powertrain system
  controller for real-time, self-learning control based on individual operating
  style},'' \emph{US Patent 8,612,107}, Dec 2013.

\bibitem{Park2011}
S.~Park, A.~Malikopoulos, M.~Kokkolaras, and D.~Jung, ``{Thermal management
  system modelling and component sizing for heavy duty series hybrid electric
  vehicles},'' \emph{International Journal of Heavy Vehicle Systems}, vol.~18,
  no.~3, pp. 272--287, 2011.

\bibitem{Malikopoulos2011}
A.~Malikopoulos and D.~Smith, ``{An optimization model for plug-in hybrid
  electric vehicles},'' in \emph{Proceedings of the 2011 Technical Conference
  of the ASME Internal Combustion Engine Division}, 2011.

\bibitem{Rios-Torres2016a}
J.~Rios-Torres and A.~Malikopoulos, ``{An overview of driver feedback systems
  for efficiency and safety},'' in \emph{IEEE Conference on Intelligent
  Transportation Systems, Proceedings, ITSC}, 2016, pp. 667--674.

\bibitem{Malikopoulos2013a}
A.~A. Malikopoulos and J.~P. Aguilar, ``{An Optimization Framework for Driver
  Feedback Systems},'' \emph{IEEE Transactions on Intelligent Transportation
  Systems}, vol.~14, no.~2, pp. 955--964, 2013.

\bibitem{Malikopoulos2016}
A.~Malikopoulos, ``{A duality framework for stochastic optimal control of
  complex systems},'' \emph{IEEE Transactions on Automatic Control}, vol.~61,
  no.~10, pp. 2756--2765, 2016.

\bibitem{Malikopoulos2013b}
------, ``{Impact of component sizing in plug-in hybrid electric vehicles for
  energy resource and greenhouse emissions reduction},'' \emph{Journal of
  Energy Resources Technology, Transactions of the ASME}, vol. 135, no.~4, p.
  041201, 2013.

\bibitem{Malikopoulos2012}
A.~Malikopoulos and J.~Aguilar, ``{Optimization of driving styles for fuel
  economy improvement},'' in \emph{IEEE Conference on Intelligent
  Transportation Systems, Proceedings, ITSC}, 2012, pp. 194--199.

\bibitem{Malikopoulos2013}
A.~Malikopoulos, ``{Stochastic optimal control for series hybrid electric
  vehicles},'' in \emph{Proceedings of the American Control Conference}, 2013,
  pp. 1191--1196.

\bibitem{Malikopoulos2014}
------, ``{Online identification of power required for self-sustainability of
  the battery in hybrid electric vehicles},'' in \emph{ASME 2014 Internal
  Combustion Engine Division Fall Technical Conference, ICEF 2014}, vol.~2,
  2014.

\bibitem{Malikopoulos2015a}
------, ``{Driver feedback for fuel efficiency},'' \emph{US Patent App.
  14/323,875}, mar 2015.

\bibitem{Malikopoulos2014a}
------, ``{Supervisory power management control algorithms for hybrid electric
  vehicles: A survey},'' \emph{IEEE Transactions on Intelligent Transportation
  Systems}, vol.~15, no.~5, pp. 1869--1885, 2014.

\bibitem{Shaltout2014}
M.~Shaltout, A.~Malikopoulos, S.~Pannala, and D.~Chen, ``{Multi-disciplinary
  decision making and optimization for hybrid electric propulsion systems},''
  in \emph{2014 IEEE International Electric Vehicle Conference, IEVC 2014},
  2014.

\bibitem{Pourazarm2014}
S.~Pourazarm, C.~G. Cassandras, and A.~Malikopoulos, ``{Optimal routing of
  electric vehicles in networks with charging nodes: A dynamic programming
  approach},'' in \emph{2014 IEEE International Electric Vehicle Conference
  (IEVC)}.\hskip 1em plus 0.5em minus 0.4em\relax IEEE, 2014, pp. 1--7.

\bibitem{Shaltout2015b}
M.~Shaltout, A.~Malikopoulos, S.~Pannala, and D.~Chen, ``{A Consumer-Oriented
  Control Framework for Performance Analysis in Hybrid Electric Vehicles},''
  \emph{IEEE Transactions on Control Systems Technology}, vol.~23, no.~4, pp.
  1451--1464, 2015.

\bibitem{Malikopoulos2015d}
A.~A. Malikopoulos, ``{Pareto Efficient Policy for Supervisory Power Management
  Control},'' in \emph{Proceedings of the 2015 IEEE 18th International
  Conference on Intelligent Transportation Systems}, 2015, pp. 2443--2448.

\bibitem{Malikopoulos2016b}
------, ``{A Multiobjective Optimization Framework for Online Stochastic
  Optimal Control in Hybrid Electric Vehicles},'' \emph{IEEE Transactions on
  Control Systems Technology}, vol.~24, no.~2, pp. 440--450, 2016.

\bibitem{Sharma2016}
I.~Sharma, J.~Dong, A.~Malikopoulos, M.~Street, J.~Ostrowski, T.~Kuruganti, and
  R.~Jackson, ``{A modeling framework for optimal energy management of a
  residential building},'' \emph{Energy and Buildings}, vol. 130, pp. 55--63,
  2016.

\bibitem{Dong2017a}
J.~Dong, T.~Kuruganti, A.~Malikopoulos, S.~Djouadi, and L.~Wang, ``{Home energy
  management based on optimal production control scheduling with unknown regime
  switching},'' in \emph{Proceedings of the American Control Conference}, 2017,
  pp. 2054--2059.

\bibitem{Dong2016}
J.~Dong, A.~Malikopoulos, S.~Djouadi, and T.~Kuruganti, ``{Application of
  optimal production control theory for home energy management in a micro
  grid},'' in \emph{Proceedings of the American Control Conference}, vol.
  2016-July, 2016, pp. 5014--5019.

\bibitem{chremos2020MobilityMarket}
I.~V. Chremos and A.~A. Malikopoulos, ``Socioeconomic impact of emerging
  mobility markets and implementation strategies,'' in \emph{AI-enabled
  Technologies for Autonomous and Connected Vehicles}, I.~Kolmanovsky,
  Y.~Murphey, and P.~Watta, Eds.\hskip 1em plus 0.5em minus 0.4em\relax
  Springer, 2023.

\bibitem{chremos2021MobilityGame}
------, ``An analytical study of a two-sided mobility game,'' in \emph{2022
  American Control Conference (ACC)}, 2022, pp. 1254--1259.

\bibitem{chremos2020SharedMobility}
------, ``Design and stability analysis of a shared mobility market,'' in
  \emph{2021 European Control Conference (ECC)}, 2021, pp. 375--380.

\bibitem{chalaki2021Reseq}
B.~Chalaki and A.~A. Malikopoulos, ``A priority-aware replanning and
  resequencing framework for coordination of connected and automated
  vehicles,'' \emph{IEEE Control Systems Letters}, vol.~6, pp. 1772--1777,
  2022.

\bibitem{Ray2021DigitalCity}
R.~M. Zayas, L.~E. Beaver, B.~Chalaki, H.~Bang, and A.~A. Malikopoulos, ``A
  digital smart city for emerging mobility systems,'' in \emph{2022 2nd Annual
  International Conference on Digital Twins and Parallel Intelligence
  (DTPI)}.\hskip 1em plus 0.5em minus 0.4em\relax IEEE, 2022, pp. 1--6.

\bibitem{mahbub2021_platoonMixed}
A.~M.~I. Mahbub and A.~A. Malikopoulos, ``{A Platoon Formation Framework in a
  Mixed Traffic Environment},'' \emph{IEEE Control Systems Letters (LCSS)},
  vol.~6, pp. 1370--1375, 2021.

\bibitem{Sumanth2021}
S.~Krishna Sumanth~Nakka, B.~Chalaki, and A.~A. Malikopoulos, ``A multi-agent
  deep reinforcement learning coordination framework for connected and
  automated vehicles at merging roadways,'' in \emph{2022 American Control
  Conference (ACC)}, 2022, pp. 3297--3302.

\bibitem{mahbub2022_ifac}
A.~M.~I. {Mahbub}, V.-A. {Le}, and A.~A. {Malikopoulos}, ``Safety-aware and
  data-driven predictive control for connected automated vehicles at a mixed
  traffic signalized intersection,'' in \emph{10th IFAC International Symposium
  on Advances in Automotive Control}.\hskip 1em plus 0.5em minus 0.4em\relax
  IFAC, 2022, pp. 51--56.

\bibitem{Malikopoulos2019CDC}
A.~A. Malikopoulos and L.~Zhao, ``Optimal path planning for connected and
  automated vehicles at urban intersections,'' in \emph{Proceedings of the 58th
  IEEE Conference on Decision and Control, 2019}.\hskip 1em plus 0.5em minus
  0.4em\relax IEEE, 2019, pp. 1261--1266.

\bibitem{Mahbub2019ACC}
A.~M.~I. Mahbub, L.~Zhao, D.~Assanis, and A.~A. Malikopoulos, ``{Energy-Optimal
  Coordination of Connected and Automated Vehicles at Multiple
  Intersections},'' in \emph{Proceedings of 2019 American Control Conference},
  2019, pp. 2664--2669.

\bibitem{malikopoulos2019ACC}
A.~A. Malikopoulos and L.~Zhao, ``A closed-form analytical solution for optimal
  coordination of connected and automated vehicles,'' in \emph{2019 American
  Control Conference (ACC)}.\hskip 1em plus 0.5em minus 0.4em\relax IEEE, 2019,
  pp. 3599--3604.

\bibitem{Malikopoulos2020}
A.~A. Malikopoulos, L.~E. Beaver, and I.~V. Chremos, ``Optimal time trajectory
  and coordination for connected and automated vehicles,'' \emph{Automatica},
  vol. 125, no. 109469, 2021.

\bibitem{mahbub2020sae-1}
A.~M.~I. Mahbub, V.~Karri, D.~Parikh, S.~Jade, and A.~A. Malikopoulos, ``A
  decentralized time- and energy-optimal control framework for connected
  automated vehicles: From simulation to field test,'' in \emph{SAE Technical
  Paper 2020-01-0579}.\hskip 1em plus 0.5em minus 0.4em\relax SAE
  International, 2020.

\bibitem{mahbub2020ACC-2}
A.~M.~I. Mahbub, A.~A. Malikopoulos, and L.~Zhao, ``Impact of connected and
  automated vehicles in a corridor,'' in \emph{Proceedings of 2020 American
  Control Conference, 2020}.\hskip 1em plus 0.5em minus 0.4em\relax IEEE, 2020,
  pp. 1185--1190.

\bibitem{chalaki2020hysteretic}
B.~Chalaki and A.~A. Malikopoulos, ``A hysteretic q-learning coordination
  framework for emerging mobility systems in smart cities,'' in \emph{2021
  European Control Conferences (ECC)}, 2021, pp. 17--22.

\bibitem{Zhao2018CTA}
L.~Zhao, A.~A. Malikopoulos, and J.~Rios-Torres, ``Optimal control of connected
  and automated vehicles at roundabouts: An investigation in a mixed-traffic
  environment,'' in \emph{15th IFAC Symposium on Control in Transportation
  Systems}, 2018, pp. 73--78.

\bibitem{tzortzoglou2023approach}
F.~N. Tzortzoglou, D.~Theodosis, and A.~Malikopoulos, ``An approach for
  optimizing acceleration in connected and automated vehicles,'' \emph{103th
  Transportation Research Board Annual Meeting (TRB)}, 2023.

\bibitem{chalaki2021CSM}
B.~Chalaki, L.~E. Beaver, A.~M.~I. Mahbub, H.~Bang, and A.~A. Malikopoulos, ``A
  research and educational robotic testbed for real-time control of emerging
  mobility systems: From theory to scaled experiments,'' \emph{IEEE Control
  Systems}, vol.~42, no.~6, pp. 20--34, 2022.

\bibitem{mahbub2020Automatica-2}
A.~M.~I. Mahbub and A.~A. Malikopoulos, ``{Conditions to Provable System-Wide
  Optimal Coordination of Connected and Automated Vehicles},''
  \emph{Automatica}, vol. 131, no. 109751, 2021.

\bibitem{jang2019simulation}
K.~Jang, E.~Vinitsky, B.~Chalaki, B.~Remer, L.~Beaver, A.~A. Malikopoulos, and
  A.~Bayen, ``Simulation to scaled city: zero-shot policy transfer for traffic
  control via autonomous vehicles,'' in \emph{Proceedings of the 10th ACM/IEEE
  International Conference on Cyber-Physical Systems}, 2019, pp. 291--300.

\bibitem{mahbub2020decentralized}
A.~I. Mahbub, A.~A. Malikopoulos, and L.~Zhao, ``Decentralized optimal
  coordination of connected and automated vehicles for multiple traffic
  scenarios,'' \emph{Automatica}, vol. 117, no. 108958, 2020.

\bibitem{Zhao2019CCTA-1}
L.~Zhao, A.~M.~I. Mahbub, and A.~A. Malikopoulos, ``Optimal vehicle dynamics
  and powertrain control for connected and automated vehicles,''
  \emph{Proceedings of 2019 IEEE Conference on Control Technology and
  Applications (CCTA)}, pp. 33--38, 2019.

\bibitem{mahbub2023_automatica}
A.~I. Mahbub, V.-A. Le, and A.~A. Malikopoulos, ``A safety-prioritized receding
  horizon control framework for platoon formation in a mixed traffic
  environment,'' \emph{Automatica}, vol. 155, p. 111115, 2023.

\bibitem{Connor2020ImpactConnectivity}
W.~D. {Connor}, Y.~{Wang}, A.~A. Malikopoulos, S.~G. {Advani}, and A.~K.
  {Prasad}, ``Impact of connectivity on energy consumption and battery life for
  electric vehicles,'' \emph{IEEE Transactions on Intelligent Vehicles},
  vol.~6, no.~1, pp. 14--23, 2020.

\bibitem{mahbub2022ACC}
A.~M.~I. {Mahbub} and A.~A. {Malikopoulos}, ``{Platoon Formation in a Mixed
  Traffic Environment: A Model-Agnostic Optimal Control Approach},''
  \emph{Proceedings of 2022 American Control Conference (ACC)}, pp. 4746--4751,
  2022.

\bibitem{Le2022CDC}
V.-A. Le and A.~A. Malikopoulos, ``{A Cooperative Optimal Control Framework for
  Connected and Automated Vehicles in Mixed Traffic Using Social Value
  Orientation},'' in \emph{2022 61th IEEE Conference on Decision and Control
  (CDC)}, 2022, pp. 6272--6277.

\bibitem{zhao2019enhanced}
L.~Zhao and A.~A. Malikopoulos, ``Enhanced mobility with connectivity and
  automation: A review of shared autonomous vehicle systems,'' \emph{IEEE
  Intelligent Transportation Systems Magazine}, vol.~14, no.~1, pp. 87--102,
  2022.

\bibitem{tzortzoglou2024feasibility}
F.~N. Tzortzoglou, L.~E. Beaver, and A.~A. Malikopoulos, ``A feasibility
  analysis at signal-free intersections,'' 2024 (in review).

\bibitem{Le2023CDC}
V.-A. Le, H.~M. Wang, G.~Orosz, and A.~A. Malikopoulos, ``{Coordination for
  Connected Automated Vehicles at Merging Roadways in Mixed Traffic
  Environment},'' in \emph{2023 62nd IEEE Conference on Decision and Control
  (CDC)}.\hskip 1em plus 0.5em minus 0.4em\relax IEEE, 2023, pp. 4150--4155.

\bibitem{Malikopoulos2018}
A.~A. Malikopoulos, C.~G. Cassandras, and Y.~J. Zhang, ``{A decentralized
  energy-optimal control framework for connected automated vehicles at
  signal-free intersections},'' \emph{Automatica}, vol.~93, pp. 244--256, 2018.

\bibitem{K122021AExperiment}
M.~Ratnagiri, C.~O'Dwyer, L.~E. Beaver, B.~Chalaki, H.~Bang, and A.~A.
  Malikopoulos, ``A scalable last-mile delivery service: From simulation to
  scaled experiment,'' in \emph{2022 25th International Conference on
  Intelligent Transportation Systems (ITSC)}.\hskip 1em plus 0.5em minus
  0.4em\relax IEEE, 2022, pp. 4163--4168.

\bibitem{Malikopoulos2017}
A.~A. Malikopoulos, C.~G. Cassandras, and Y.~Zhang, ``A decentralized
  energy-optimal control framework for connected automated vehicles at
  signal-free intersections,'' \emph{Automatica}, vol.~93, pp. 244 -- 256,
  2018.

\bibitem{Le2023Stochastic}
V.-A. Le, B.~Chalaki, F.~N. Tzortzoglou, and A.~A. Malikopoulos, ``{Stochastic
  Time-Optimal Trajectory Planning for Connected and Automated Vehicles in
  Mixed-Traffic Merging Scenarios},'' \emph{arXiv preprint arXiv:2311.00126},
  2023.

\bibitem{Malikopoulos2018b}
A.~A. Malikopoulos, S.~Hong, B.~B. Park, J.~Lee, and S.~Ryu, ``{Optimal control
  for speed harmonization of automated vehicles},'' \emph{IEEE Transactions on
  Intelligent Transportation Systems}, vol.~20, no.~7, pp. 2405--2417, 2019.

\bibitem{mahbub2020sae-2}
A.~M.~I. Mahbub and A.~A. Malikopoulos, ``Concurrent optimization of vehicle
  dynamics and powertrain operation using connectivity and automation,'' in
  \emph{SAE Technical Paper 2020-01-0580}.\hskip 1em plus 0.5em minus
  0.4em\relax SAE International, 2020.

\bibitem{le2024controller}
V.-A. Le and A.~A. Malikopoulos, ``{Controller Adaptation via Learning
  Solutions of Contextual Bayesian Optimization},'' \emph{arXiv preprint
  arXiv:2403.04881}, 2024.

\bibitem{Rios-Torres2015}
J.~Rios-Torres, A.~Malikopoulos, and P.~Pisu, ``{Online Optimal Control of
  Connected Vehicles for Efficient Traffic Flow at Merging Roads},'' in
  \emph{IEEE Conference on Intelligent Transportation Systems, Proceedings,
  ITSC}, 2015, pp. 2432--2437.

\bibitem{Rios-Torres2017b}
J.~Rios-Torres and A.~A. Malikopoulos, ``{Automated and Cooperative Vehicle
  Merging at Highway On-Ramps},'' \emph{IEEE Transactions on Intelligent
  Transportation Systems}, vol.~18, no.~4, pp. 780--789, 2017.

\bibitem{Rios-Torres2017a}
------, ``{Impact of connected and automated vehicles on traffic flow},'' in
  \emph{2017 IEEE 20th International Conference on Intelligent Transportation
  Systems (ITSC)}, 2017, pp. 1--6.

\bibitem{Rios-Torres2017}
J.~Rios-Torres and A.~Malikopoulos, ``{A Survey on the Coordination of
  Connected and Automated Vehicles at Intersections and Merging at Highway
  On-Ramps},'' \emph{IEEE Transactions on Intelligent Transportation Systems},
  vol.~18, no.~5, pp. 1066--1077, 2017.

\bibitem{Rios-Torres2016b}
------, ``{Energy impact of different penetrations of connected and automated
  vehicles: A preliminary assessment},'' in \emph{Proceedings of the 9th ACM
  SIGSPATIAL International Workshop on Computational Transportation Science,
  IWCTS 2016}, 2016.

\bibitem{Zhao2018ITSC}
L.~Zhao and A.~A. Malikopoulos, ``Decentralized optimal control of connected
  and automated vehicles in a corridor,'' in \emph{2018 21st International
  Conference on Intelligent Transportation Systems (ITSC)}, Nov 2018, pp.
  1252--1257.

\bibitem{tzortzoglou2023performance}
F.~N. Tzortzoglou, D.~Theodosis, A.~Dave, and A.~Malikopoulos,
  ``Performance-sensitive potential functions for efficient flow of connected
  and automated vehicles,'' \emph{Accepted for the proceeding of the American
  Control Conference 2024 (ACC)}, 2023.

\bibitem{zhao2019CCTA-2}
L.~Zhao, A.~A. Malikopoulos, and J.~Rios-Torres, ``On the traffic impacts of
  optimally controlled connected and automated vehicles,'' in \emph{Proceedings
  of 2019 IEEE Conference on Control Technology and Applications (CCTA)}.\hskip
  1em plus 0.5em minus 0.4em\relax IEEE, 2019, pp. 882--887.

\bibitem{Rios-Torres2018}
J.~Rios-Torres and A.~A. Malikopoulos, ``{Impact of partial penetrations of
  connected and automated vehicles on fuel consumption and traffic flow},''
  \emph{IEEE Transactions on Intelligent Vehicles}, vol.~3, no.~4, pp.
  453--462, 2018.

\bibitem{Stager}
A.~Stager, L.~Bhan, A.~Malikopoulos, and L.~Zhao, ``{A Scaled Smart City for
  Experimental Validation of Connected and Automated Vehicles},'' in \emph{15th
  IFAC Symposium on Control in Transportation Systems: CTS 2018}, pp. 130--135.

\bibitem{Zhao2018}
L.~Zhao, A.~Malikopoulos, and J.~Rios-Torres, ``{Optimal control of connected
  and automated vehicles at roundabouts: An investigation in a mixed-traffic
  environment},'' \emph{IFAC-PapersOnLine}, vol.~51, no.~9, pp. 73--78, 2018.

\bibitem{Zhang2017a}
Y.~Zhang, C.~Cassandras, and A.~Malikopoulos, ``{Optimal control of Connected
  Automated Vehicles at urban traffic intersections: A feasibility enforcement
  analysis},'' in \emph{Proceedings of the American Control Conference}, 2017,
  pp. 3548--3553.

\bibitem{Zhang}
Y.~Zhang, A.~Malikopoulos, and C.~G. Cassandras, ``{Decentralized Optimal
  Control for Connected and Automated Vehicles at Intersections Including Left
  and Right Turns},'' in \emph{56th IEEE Conference on Decision and Control},
  2017, pp. 4428--4433.

\bibitem{Zhao2018a}
L.~Zhao and A.~A. Malikopoulos, ``{Decentralized optimal control of connected
  and automated vehicles in a corridor},'' in \emph{2018 21st International
  Conference on Intelligent Transportation Systems (ITSC)}, 2018, pp.
  1252--1257.

\bibitem{assanis2018ITSC}
D.~Assanis, L.~Zhao, and A.~A. Malikopoulos, ``Characterization of the new
  class of driving cycles for connected and automated vehicles,'' in \emph{2018
  21st International Conference on Intelligent Transportation Systems
  (ITSC)}.\hskip 1em plus 0.5em minus 0.4em\relax IEEE, 2018, pp. 3668--3673.

\bibitem{Beaver2020DemonstrationCity}
L.~E. Beaver, B.~Chalaki, A.~M. Mahbub, L.~Zhao, R.~Zayas, and A.~A.
  Malikopoulos, ``{Demonstration of a Time-Efficient Mobility System Using a
  Scaled Smart City},'' \emph{Vehicle System Dynamics}, vol.~58, no.~5, pp.
  787--804, 2020.

\bibitem{chalaki2019optimal}
B.~Chalaki and A.~A. Malikopoulos, ``An optimal coordination framework for
  connected and automated vehicles in two interconnected intersections,'' in
  \emph{2019 IEEE Conference on Control Technology and Applications
  (CCTA)}.\hskip 1em plus 0.5em minus 0.4em\relax IEEE, 2019, pp. 888--893.

\bibitem{chalaki2020experimental}
B.~Chalaki, L.~E. Beaver, and A.~A. Malikopoulos, ``Experimental validation of
  a real-time optimal controller for coordination of cavs in a multi-lane
  roundabout,'' in \emph{31st IEEE Intelligent Vehicles Symposium (IV)}, 2020,
  pp. 504--509.

\bibitem{chalaki2020ICCA}
B.~Chalaki, L.~E. Beaver, B.~Remer, K.~Jang, E.~Vinitsky, A.~Bayen, and A.~A.
  Malikopoulos, ``Zero-shot autonomous vehicle policy transfer: From simulation
  to real-world via adversarial learning,'' in \emph{IEEE 16th International
  Conference on Control \& Automation (ICCA)}, 2020, pp. 35--40.

\bibitem{Mahbub2020ACC-1}
A.~M.~I. Mahbub and A.~A. Malikopoulos, ``Conditions for state and control
  constraint activation in coordination of connected and automated vehicles,''
  \emph{Proceedings of 2020 American Control Conference}, pp. 436--441, 2020.

\bibitem{chalaki2020TCST}
B.~Chalaki and A.~A. Malikopoulos, ``Optimal control of connected and automated
  vehicles at multiple adjacent intersections,'' \emph{IEEE Transactions on
  Control Systems Technology}, vol.~30, no.~3, pp. 972--984, 2022.

\bibitem{Nishanth2023AISmerging}
N.~Venkatesh, V.-A. Le, A.~Dave, and A.~A. Malikopoulos, ``Connected and
  automated vehicles in mixed-traffic: Learning human driver behavior for
  effective on-ramp merging,'' in \emph{Proceedings of the 62nd IEEE Conference
  on Decision and Control (CDC)}.\hskip 1em plus 0.5em minus 0.4em\relax IEEE,
  2023, pp. 92--97.

\bibitem{chalaki2020TITS}
B.~Chalaki and A.~A. Malikopoulos, ``Time-optimal coordination for connected
  and automated vehicles at adjacent intersections,'' \emph{IEEE Transactions
  on Intelligent Transportation Systems}, 2021.

\bibitem{bang2024confidence}
H.~Bang, A.~Dave, and A.~A. Malikopoulos, ``Safe merging in mixed traffic with
  confidence,'' 2024 (in review).

\bibitem{Bang2023flowbased}
H.~Bang and A.~A. Malikopoulos, ``A hierarchical approach to optimal flow-based
  routing and coordination of connected and automated vehicles,'' in
  \emph{Proceedings of the 62nd IEEE Conference on Decision and Control (CDC)},
  2023, pp. 7100--7105.

\bibitem{bang2021AEMoD}
------, ``Congestion-aware routing, rebalancing, and charging scheduling for
  electric autonomous mobility-on-demand system,'' in \emph{Proceedings of 2022
  American Control Conference (ACC)}, 2022, pp. 3152--3157.

\bibitem{bang2023exploring}
H.~Bang, J.~Holden, A.~Mahbub, J.~Gonder, and A.~A. Malikopoulos, ``Exploring
  microsimulation process for energy impact evaluation of connected and
  automated vehicles,'' in \emph{Transportation Research Board 102nd Annual
  Meeting}, 2023.

\bibitem{bang2024cts}
H.~Bang, A.~Dave, F.~Tzortzoglou, and A.~A. Malikopoulos, ``A mobility equity
  metric for multi-modal intelligent transportation systems,'' 2024 (to
  apprear).

\bibitem{Chremos2020MechanismDesign}
I.~V. Chremos and A.~A. Malikopoulos, ``Social resource allocation in a
  mobility system with connected and automated vehicles: A mechanism design
  problem,'' in \emph{Proceedings of the 59th IEEE Conference on Decision and
  Control (CDC), 2020}, 2020, pp. 2642--2647.

\bibitem{Bang2022rerouting}
H.~Bang and A.~A. Malikopoulos, ``{{Re-Routing Strategy of Connected and
  Automated Vehicles Considering Coordination at Intersections}},'' in
  \emph{Proceedings of 2023 American Control Conference (ACC)}, 2023, pp.
  4419--4424.

\bibitem{Bang2022combined}
H.~Bang, B.~Chalaki, and A.~A. Malikopoulos, ``{Combined Optimal Routing and
  Coordination of Connected and Automated Vehicles},'' \emph{IEEE Control
  Systems Letters}, vol.~6, pp. 2749--2754, 2022.

\bibitem{chremos2023AtomicRouting}
I.~V. Chremos, H.~Bang, A.~Dave, V.-A. Le, and A.~A. Malikopoulos, ``Modeling
  travel behavior in mobility systems with an atomic routing game and prospect
  theory,'' \emph{arXiv:2303.17790}, 2023.

\bibitem{bang2023optimal}
H.~Bang and A.~A. Malikopoulos, ``Optimal trajectory planning meets
  network-level routing: Integrated control framework for emerging mobility
  systems,'' 2023 (in review, arXiv:2311.13193).

\bibitem{bang2024emergingequity}
H.~Bang, A.~Dave, F.~Tzortzoglou, S.~Wang, and A.~A. Malikopoulos, ``On
  mobility equity and the promise of emerging transportation systems,''
  \emph{IEEE Transactions on Intelligent Transportation Systems}, 2024 (in
  review).

\bibitem{Bang2023mem}
H.~Bang, A.~Dave, and A.~A. Malikopoulos, ``{Routing in Mixed Transportation
  Systems for Mobility Equity},'' \emph{Proceedings of the 2023 American
  Control Conference}, 2023 (to appear, arXiv:2309.03981).

\bibitem{dave2024airecommend}
A.~Dave, H.~Bang, and A.~A. Malikopoulos, ``A framework for effective ai
  recommendations in cyber-physical-human systems,'' \emph{IEEE Control Systems
  Letters}, 2024 (accepted).

\bibitem{faros2023adherence}
I.~Faros, A.~Dave, and A.~A. Malikopoulos, ``A q-learning approach for
  adherence-aware recommendations,'' \emph{IEEE Control Systems Letters},
  vol.~7, pp. 3645--3650, 2023.

\bibitem{Malikopoulos2024}
A.~A. Malikopoulos, ``Combining learning and control in linear systems,'' in
  \emph{22nd European Control Conference (ECC)}, 2024 (to appear) (arxiv:
  2310.14409).

\bibitem{Malikopoulos2021}
------, ``On team decision problems with nonclassical information structures,''
  \emph{IEEE Transactions on Automatic Control}, vol.~68, no.~7, pp.
  3915--3930, 2023.

\bibitem{Dave2021minimax}
A.~Dave, N.~Venkatesh, and A.~A. Malikopoulos, ``On decentralized minimax
  control with nested subsystems,'' in \emph{2022 American Control Conference
  (ACC)}.\hskip 1em plus 0.5em minus 0.4em\relax IEEE, 2022, pp. 3437--3444.

\bibitem{dave2019decentralized}
A.~Dave and A.~A. Malikopoulos, ``Decentralized stochastic control in partially
  nested information structures,'' \emph{IFAC-PapersOnLine}, vol.~52, no.~20,
  pp. 97--102, 2019.

\bibitem{Dave2020}
------, ``{The Prescription Approach for Decentralized Stochastic Control with
  Word-of-Mouth Communication},'' \emph{arXiv preprint, arXiv:1907.12125},
  2021.

\bibitem{Dave2021nestedaccess}
A.~Dave, N.~Venkatesh, and A.~A. Malikopoulos, ``Decentralized control of two
  agents with nested accessible information,'' in \emph{2022 American Control
  Conference (ACC)}.\hskip 1em plus 0.5em minus 0.4em\relax IEEE, 2022, pp.
  3423--3430.

\bibitem{Dave2020a}
A.~Dave and A.~A. Malikopoulos, ``Structural results for decentralized
  stochastic control with a word-of-mouth communication,'' in \emph{2020
  American Control Conference (ACC)}.\hskip 1em plus 0.5em minus 0.4em\relax
  IEEE, 2020, pp. 2796--2801.

\bibitem{Dave2021a}
------, ``A dynamic program for a team of two agents with nested information,''
  in \emph{2021 IEEE Conference on Decision and Control (CDC)}.\hskip 1em plus
  0.5em minus 0.4em\relax IEEE, 2021, pp. 3768--3773.

\bibitem{Dave2023infhorizon}
A.~Dave, I.~Faros, N.~Venkatesh, and A.~A. Malikopoulos, ``Worst-case control
  and learning using partial observations over an infinite time horizon,'' in
  \emph{Proceedings of the 62nd IEEE Conference on Decision and Control
  (CDC)}.\hskip 1em plus 0.5em minus 0.4em\relax IEEE, 2023, pp. 6014--6019.

\bibitem{venkatesh2023stochastic}
N.~Venkatesh, A.~Dave, I.~Faros, and A.~A. Malikopoulos, ``Stochastic control
  with distributionally robust constraints for cyber-physical systems
  vulnerable to attacks,'' \emph{European Journal of Control}, 2024
  (accepted)(arXiv:2311.03666).

\bibitem{dave2022additive}
A.~Dave, N.~Venkatesh, and A.~A. Malikopoulos, ``On robust control of partially
  observed uncertain systems with additive costs,'' in \emph{Proceedings of the
  2023 American Control Conference (ACC)}, 2023, pp. 4639--4644.

\bibitem{Dave2023approximate}
------, ``{Approximate Information States for Worst-Case Control and Learning
  in Uncertain Systems},'' \emph{arXiv:2301.05089 (in review)}, 2023.

\bibitem{Dave2022approx}
------, ``Approximate information states for worst-case control of uncertain
  systems,'' in \emph{Proceedings of the 61th IEEE Conference on Decision and
  Control (CDC)}, 2022, pp. 4945--4950.

\bibitem{chremos2022CSMArticle}
I.~V. Chremos and A.~A. Malikopoulos, ``Mechanism design theory in control
  engineering: A tutorial and overview of applications in communication, power
  grid, transportation, and security systems,'' \emph{IEEE Control Systems},
  vol.~44, no.~1, pp. 20--45, 2024.

\bibitem{Dave2020SocialMedia}
A.~Dave, I.~V. Chremos, and A.~A. Malikopoulos, ``Social media and misleading
  information in a democracy: A mechanism design approach,'' \emph{IEEE
  Transactions on Automatic Control}, vol.~67, no.~5, pp. 2633--2639, 2022.

\bibitem{Beaver2020AnFlockingb}
L.~E. Beaver and A.~A. Malikopoulos, ``{An Overview on Optimal Flocking},''
  \emph{Annual Reviews in Control}, vol.~51, pp. 88--99, 2021.

\bibitem{Beaver2020AnAgents}
------, ``{An Energy-Optimal Framework for Assignment and Trajectory Generation
  in Teams of Autonomous Agents},'' \emph{Systems \& Control Letters}, vol.
  138, April 2020.

\bibitem{Beaver2020Energy-OptimalConstraints}
L.~E. Beaver, M.~Dorothy, C.~Kroninger, and A.~A. Malikopoulos,
  ``{Energy-Optimal Motion Planning for Agents: Barycentric Motion and
  Collision Avoidance Constraints},'' in \emph{2021 American Control
  Conference}, 2021, pp. 1037--1042.

\bibitem{bang2021energy}
H.~Bang, L.~E. Beaver, and A.~A. Malikopoulos, ``Energy-optimal goal assignment
  of multi-agent system with goal trajectories in polynomials,'' in \emph{2021
  29th Mediterranean Conference on Control and Automation (MED)}.\hskip 1em
  plus 0.5em minus 0.4em\relax IEEE, 2021, pp. 1228--1233.

\bibitem{Beaver2019AGeneration}
L.~E. Beaver and A.~A. Malikopoulos, ``{A Decentralized Control Framework for
  Energy-Optimal Goal Assignment and Trajectory Generation},'' in \emph{IEEE
  58th Conference on Decision and Control}, 2019, pp. 879--884.

\bibitem{Beaver2021Constraint-DrivenStudy}
------, ``Constraint-driven optimal control of multi-agent systems: A highway
  platooning case study,'' \emph{IEEE Control Systems Letters}, vol.~6, pp.
  1754--1759, 2022.

\bibitem{Beaver2020BeyondFlocking}
------, ``{Beyond Reynolds: A Constraint-Driven Approach to Cluster
  Flocking},'' in \emph{IEEE 59th Conference on Decision and Control}, 2020,
  pp. 208--213.

\bibitem{Beaver2020AnFlocking}
L.~E. Beaver, C.~Kroninger, and A.~A. Malikopoulos, ``{An Optimal Control
  Approach to Flocking},'' in \emph{2020 American Control Conference}, 2020,
  pp. 683--688.

\end{thebibliography}

\documentclass[letterpaper,11pt]{article}
\usepackage[letterpaper, left=1in,right=1in,top=1in,bottom=1in]{geometry}   
\usepackage{color}
\usepackage{amssymb}
\usepackage{epsfig}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{graphicx}
\usepackage{subfigure}
\usepackage{color}
\usepackage{graphicx}
\usepackage{epsfig}
\usepackage{cite}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{mathrsfs}
\usepackage{amsfonts}
\usepackage{dsfont}
\usepackage{graphicx}
\usepackage{mathrsfs}
\usepackage{hyperref}
\usepackage{hyperref}
\usepackage{titlesec}
\usepackage[T1]{fontenc}
\usepackage{ragged2e}
\justifying

\DeclareGraphicsExtensions{.pdf,.png,.jpg,.JPG}

\usepackage{lastpage}
\usepackage{fancyhdr}
\raggedright
\titleformat{\section}{
	\vspace{12pt}\scshape\raggedright\large
}{}{0em}{\textsc}

\begin{document}
	\thispagestyle{empty}
	
	\begin{tabular*}{\textwidth}{l@{\extracolsep{\fill}}r}
		\textbf{\centerline{A Note for CPS Data-driven Approaches Developed in the IDS Lab}}\vspace{5pt}\\ 
		\centerline{Andreas A. Malikopoulos}\\
        \centerline{Cornell University}\\
        \centerline{amaliko@cornell.edu}\\
	\end{tabular*}

\vspace{+5EX}

\centerline{\textbf{Abstract}}

\justify{

The rapid evolution of Cyber-Physical Systems (CPS) across various domains like mobility systems, networked control systems, sustainable manufacturing, smart power grids, and the Internet of Things  necessitates innovative solutions that merge control and learning \cite{Malikopoulos2022a}. Traditional model-based control methodologies often fail to adapt to the dynamism and complexity of modern CPS. This report outlines a comprehensive approach undertaken by the Information and Decision Science (IDS) Lab, focusing on integrating data-driven techniques with control strategies to enhance CPS performance, particularly in the context of energy efficiency and environmental impact.
CPS are intricate networks where physical and software components are deeply intertwined, operating as systems of systems. These systems are characterized by their informationally decentralized nature, posing significant challenges in optimization and control. Classical control methods depend heavily on precise models, which often do not capture the full complexity of real-world CPS. As these systems generate large volumes of real-time data, there is a growing need for control algorithms that can leverage this data effectively. The IDS Lab is at the forefront of developing such data-driven approaches for CPS.
	
	\section{\textbf{Autonomous Powertrain Systems}}
	\justify{
		
		We have conducted research on autonomous powertrain control systems capable of learning and adapting to drivers’ behaviors to optimize engine performance and fuel efficiency \cite{Malikopoulos2011b, Malikopoulos2007,Malikopoulos2007a,Malikopoulos2009b, Malikopoulos2009a,Malikopoulos2006,Malikopoulos2011a,Malikopoulos2015,Malikopoulos2015b,Malikopoulos2008b,Malikopoulos2010a,Malikopoulos2010,Malikopoulos2009e,Malikopoulos2010b, Malikopoulos2009,Malikopoulos2013d}. This interest was sparked by an article highlighting the significant impact of driving styles on vehicle performance. We have developed a theoretical framework for transforming a vehicle’s engine into an intelligent system that learns its optimal operation dynamically. By modeling the driver’s behavior as a control Markov chain, the engine learns and adapts to achieve optimal performance criteria, such as minimizing fuel consumption and emissions. This framework’s practical applications led to a US patent and subsequent developments at General Motors, where I contributed to creating self-learning control algorithms for advanced powertrain systems.

		We have also broadened our research to include stochastic optimal control and power management for hybrid-electric vehicles  and plug-in hybrid-electric vehicles  	\cite{Park2011,Malikopoulos2011,Rios-Torres2016a,Malikopoulos2013a,Malikopoulos2016,Malikopoulos2013b,Malikopoulos2012,Malikopoulos2013,Malikopoulos2014,Malikopoulos2015a,Malikopoulos2014a,Shaltout2014,Pourazarm2014, Shaltout2015b,Malikopoulos2015d,Malikopoulos2016b}. Our work involved developing algorithms for optimal energy distribution and power management in these vehicles, enhancing their efficiency and sustainability. In addition, we have investigated the integration of these vehicles with smart grids and the implications of their widespread adoption on urban infrastructure \cite{Sharma2016,Dong2017a,Dong2016}. My role as Deputy Director of the Urban Dynamics Institute at Oak Ridge National Laboratory allowed me to spearhead initiatives that examined the environmental impacts of conected and automated vehicles (CAVs) and sought to improve urban transportation systems’ sustainability and accessibility.
	}

\section{\textbf{Emerging Mobility Systems}}
\justify{
	
Emerging mobility systems \cite{chremos2020MobilityMarket,chremos2021MobilityGame,chremos2020SharedMobility} represent a new paradigm in transportation, integrating advanced technologies and control algorithms to enhance efficiency and sustainability 	\cite{chalaki2021Reseq, Ray2021DigitalCity,mahbub2021_platoonMixed,Sumanth2021,mahbub2022_ifac, Malikopoulos2019CDC, Mahbub2019ACC,  malikopoulos2019ACC, Malikopoulos2020, mahbub2020sae-1, mahbub2020ACC-2, chalaki2020hysteretic, Zhao2018CTA, tzortzoglou2023approach, chalaki2021CSM,mahbub2020Automatica-2, jang2019simulation, mahbub2020decentralized, Zhao2019CCTA-1, mahbub2023_automatica, Connor2020ImpactConnectivity, mahbub2022ACC,Le2022CDC, zhao2019enhanced, tzortzoglou2024feasibility, Le2023CDC, Malikopoulos2018,K122021AExperiment,Malikopoulos2017,Le2023Stochastic, Malikopoulos2018b,mahbub2020sae-2, le2024controller, Rios-Torres2015,Rios-Torres2017b,Rios-Torres2017a,Rios-Torres2017,Rios-Torres2016b,Zhao2018ITSC, tzortzoglou2023performance,zhao2019CCTA-2,  Rios-Torres2018,Stager,Zhao2018,Zhang2017a,Zhang,Zhao2018a, assanis2018ITSC,Beaver2020DemonstrationCity,chalaki2019optimal,chalaki2020experimental,chalaki2020ICCA,Mahbub2020ACC-1, chalaki2020TCST,Nishanth2023AISmerging,chalaki2020TITS}. These systems include CAVs, shared mobility platforms, and interconnected public transit networks. The IDS Lab's research in this area has focused on developing control algorithms for the optimal coordination of CAVs, addressing challenges such as collision avoidance, traffic flow optimization, and energy efficiency. One significant advancement was the development of a decentralized optimal control framework for CAVs at merging roadways, which significantly improves fuel efficiency and reduces emissions by optimizing vehicle trajectories in real-time.	
	
Autonomous Mobility-on-Demand (AMoD) systems \cite{bang2024confidence, Bang2023flowbased,bang2021AEMoD, bang2023exploring,bang2024cts} represent a significant evolution in urban transportation \cite{Chremos2020MechanismDesign}, offering dynamic, real-time routing \cite{Bang2022rerouting, Bang2022combined,chremos2023AtomicRouting,bang2023optimal} and scheduling of autonomous vehicles. The IDS Lab's research in this area has focused on developing algorithms for the efficient operation of electric AMoD systems, including routing optimization and charging infrastructure planning. By integrating CAV behavior into AMoD systems, we have developed hierarchical optimization frameworks that improve traffic flow and system scalability. This research addresses the technical challenges of AMoD systems while also considers the broader impacts on urban mobility, sustainability, and equity \cite{bang2024emergingequity,bang2024cts,Bang2023mem}.
}

\section{\textbf{Human Involvement in Cyber-Physical Systems}}
\justify{
Considering human factors in the design and operation of CPS is crucial, particularly in socio-technical systems where human behavior significantly impacts system performance. The IDS Lab's in this area research has explored the interactions between AI-driven systems and human users, developing algorithms that adapt to human behaviors to optimize system performance \cite{dave2024airecommend,faros2023adherence}. Additionally, I have proposed novel metrics for assessing mobility equity in transportation networks, ensuring that advancements in CPS contribute to social equity and inclusivity.

}

\section{\textbf{Team Theory}}
\justify{

Team theory offers a valuable framework for addressing decentralized control problems involving multiple agents with shared objectives but different information. The IDS Lab's research in this domain has focused on sequential dynamic team decision problems, providing structural results and a dynamic programming decomposition for such problems \cite{Malikopoulos2024,Malikopoulos2022a,Malikopoulos2021,Dave2021minimax,dave2019decentralized,Dave2020,Dave2021nestedaccess,Dave2020a,Dave2021a}. These contributions have important implications for the design and control of informationally decentralized systems like CPS, enabling more efficient and robust coordination among system components \cite{Nishanth2023AISmerging,Dave2023infhorizon,venkatesh2023stochastic,dave2022additive,Dave2023approximate, Dave2022approx}.

}

\section{\textbf{Mechanism Design Theory}}
\justify{

Mechanism design theory, a subfield of game theory, provides a structured approach to designing rules and protocols that lead to desired outcomes, even when participants have private information and individual incentives. The IDS Lab's research in this area has focused on applying mechanism design to CPS, particularly in contexts where agents need to make decisions that are socially optimal \cite{chremos2022CSMArticle,Dave2020SocialMedia}.
One major application of mechanism design in my research has been in the development of incentive schemes for energy-efficient behavior in mobility markets \cite{chremos2023AtomicRouting}. By designing mechanisms that align individual incentives with collective goals, we can encourage behaviors that enhance system-wide efficiency and sustainability.

}

\section{\textbf{Multi-Agent System}}
\justify{

Multi-agent systems  play a critical role in the landscape of CPS, where numerous autonomous entities interact and collaborate to achieve complex tasks \cite{Beaver2020AnFlockingb}. The IDS Lab's research in this area has focused on developing robust coordination algorithms that ensure seamless interaction among agents, which may include vehicles, drones, or other autonomous systems  \cite{Beaver2020AnAgents,Beaver2020Energy-OptimalConstraints,bang2021energy,Beaver2019AGeneration,Beaver2021Constraint-DrivenStudy,Beaver2020BeyondFlocking,Beaver2020AnFlocking}. These algorithms address critical challenges such as task allocation, resource sharing, conflict resolution, and cooperative decision-making. One notable project involved designing a decentralized control framework for multi-agent systems, enabling agents to operate with limited information about the global state while still achieving collective goals. This approach leverages principles of game theory and distributed optimization to ensure that agents can make informed decisions based on local observations and interactions. In the context of emerging mobility systems, MAS research has been pivotal in optimizing the behavior of connected and automated vehicles within mixed traffic environments. By simulating various scenarios, we have developed strategies that enhance traffic flow, reduce congestion, and improve safety. These strategies involve real-time data exchange among vehicles and infrastructure, enabling adaptive responses to dynamic traffic conditions. 
Furthermore, my research has explored the integration of MAS with machine learning techniques, particularly reinforcement learning, to enable agents to learn and adapt to their environment over time. This combination of MAS and learning algorithms has shown promise in applications ranging from urban air mobility to automated logistics and smart manufacturing.

}

\newpage

\bibliographystyle{IEEEtran}
\bibliography{references1,IDS_Publications_06132024,TAC_Ref_structure,TAC_Ref_Andreas}

\end{document}